Solve using differentiation rule


y = 2x ^ 2 * sqrt(2 - x)

1
Expert's answer
2022-05-18T09:47:22-0400
"y'=(2x^2\sqrt{2-x})'=4x\sqrt{2-x}+\dfrac{2x^2}{2\sqrt{2-x}(-1)}"

"=4x\sqrt{2-x}-\dfrac{x^2}{\sqrt{2-x}}"

"=\dfrac{4x(2-x)-x^2}{2\sqrt{2-x}}=\dfrac{8x-5x^2}{2\sqrt{2-x}}"